# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 1.1. Background and motivation
Let $I$ be a pro-$p$ Iwahori subgroup of $G$ and let ${\mathcal{H}}=k[I\backslash G/I]$ be the corresponding pro-$p$ Iwahori-Hecke algebra of $G$, i.e. the convolution algebra of $I$-biinvariant compactly supported functions on $G$. Write ${\operatorname{Mod}}({\mathcal{H}})$ for the category of right ${\mathcal{H}}$-modules. Then taking $I$-invariants defines a functor $U:{\operatorname{Rep}}_{k}^{\infty}(G)\to{\operatorname{Mod}}({\mathcal{H}})$ which sends non-zero representations to non-zero modules. This functor is expected to provide a strong relation between the two categories, however its behaviour is not sufficiently understood beyond $G=\mathrm{GL}_{2}(\mathbb{Q}_{p})$ and a few related cases (c.f. the work of Ollivier [24]). In particular, one of the major open problems in the theory is to determine the relationship between the isomorphism classes of irreducible admissible smooth representations of $G$ and of simple ${\mathcal{H}}$-modules.
In previous work [9] with Kohlhaase, we showed that $U$ restricts to an equivalence between an explicit full subcategory of ${\operatorname{Rep}}_{k}^{\infty}(G)$ and the full subcategory $\mathrm{GProj}({\mathcal{H}})$ of Gorenstein projective ${\mathcal{H}}$-modules. These modules are a key component in the definition of a model structure on ${\operatorname{Mod}}({\mathcal{H}})$ due to Hovey [15]. We used this in [9] to construct and study a model structure on ${\operatorname{Rep}}_{k}^{\infty}(G)$ and the derived functor $\mathbf{R}U$ between the associated homotopy categories. We generally obtained that the situation was improved at the homotopy level. This suggests that one should instead investigate the relationship between the homotopy classes of irreducible admissible smooth representations of $G$ and of simple ${\mathcal{H}}$-modules.
@@
Our main goal in this work is to treat the harder problem of determining whether two nonisomorphic simple ${\mathcal{H}}$-modules of infinite projective dimension can become isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. If $M$ denotes a standard Levi subgroup of $G$ with associated Hecke algebra ${\mathcal{H}}_{M}$ then a key ingredient in Abe’s construction of the simple modules in [1] is the parabolic induction functor ${\operatorname{Ind}}^{{\mathcal{H}}}_{{\mathcal{H}}_{M}}:{\operatorname{Mod}}({\mathcal{H}}_{M})\to{\operatorname{Mod}}({\mathcal{H}})$. We therefore first aim to study the properties of ${\operatorname{Ind}}^{{\mathcal{H}}}_{{\mathcal{H}}_{M}}$ in the language of Hovey’s Gorenstein projective model structure. This allows us to essentially reduce our question to the case of supersingular modules. In many situations of interest, we are then able to obtain a rather complete understanding of the situation (c.f. B & C below for details).
@@ 1.2. Main results
[…]
@@
This is proved in 3.13 and 3.31. Roughly speaking, part (i) of A says that the triangulated category ${\operatorname{Ho}}({\mathcal{H}})$ can be decomposed into two orthogonal thick subcategories, one equal to ${\mathcal{S}}^{G}_{\mathcal{M}}$ and the other one equal to ${\operatorname{Ho}}({\mathcal{H}})_{\mathcal{M}}$, where the latter category is embedded fully faithfully into ${\operatorname{Ho}}({\mathcal{H}})$ via either one of the functors $\ell^{G}_{\mathcal{M}}$ or $r^{G}_{\mathcal{M}}$ (c.f. 3.9 and the discussion following it for details). We point out that while the proof of (i) is mostly formal, it however relies on the fact that parabolic induction is fully faithful (c.f. 2.5), which appears to be a new result.
@@
The proof of part (ii) of A relies on Abe’s classification of the simple ${\mathcal{H}}$-modules in [1], on Koziol’s results on the projective dimension of simple modules from [19], and on a lift to the homotopy level of a computation of Abe of the image of the simple modules under the adjoints of the parabolic induction functor (c.f. 3.27). As a consequence of the latter computation, we also obtain a new proof of a result of Koziol and reduce the study of isomorphism classes of simple ${\mathcal{H}}$-modules in ${\operatorname{Ho}}({\mathcal{H}})$ to the case of simple supersingular ${\mathcal{H}}$-modules, c.f. 3.28.
Having established this, the next step is to determine when two simple supersingular ${\mathcal{H}}$-modules can become isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. Here the main difficulty is that computing the Hom spaces in ${\operatorname{Ho}}({\mathcal{H}})$ a priori requires to compute explicit Gorenstein projective replacements in ${\operatorname{Mod}}({\mathcal{H}})$. Instead, inspired by the arguments of Koziol in [19] and by the functorial resolutions of Ollivier-Schneider [25], we use Bruhat-Tits theory. If $F$ denotes a face of the Bruhat-Tits building contained in the closure of a chosen chamber $C$, then there is an associated subalgebra ${\mathcal{H}}_{F}$ of ${\mathcal{H}}$. These subalgebras together generate a Gorenstein subalgebra ${\mathcal{H}_{\operatorname{aff}}}$ of ${\mathcal{H}}$ called the affine Hecke algebra. The simple supersingular Hecke modules were classified by Vignéras, c.f. [32], and a key input in this classification are the so-called supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$. Our main insight is to consider the ‘diagonal’ functor
@@
which is right Quillen and in fact determines the model structure on ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ via a right transfer property, c.f. 5.5. This indicates that we should in principle be able to do computations in the product category $\prod_{F\subseteq\overline{C}}{\operatorname{Ho}}({\mathcal{H}}_{F})$ instead of ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$. The advantage being that since ${\mathcal{H}}_{F}$ is known to be self-injective, it follows that every ${\mathcal{H}}_{F}$-module is Gorenstein projective and hence there is no need to compute replacements in order to compute Hom spaces in ${\operatorname{Ho}}({\mathcal{H}}_{F})$.
More specifically, we have that $\Delta$ lifts to a functor ${\operatorname{Ho}}(\Delta):{\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})\to\prod_{F\subseteq\overline{C}}{\operatorname{Ho}}({\mathcal{H}}_{F})$. One major technical input in this paper is the computation of the image under ${\operatorname{Ho}}(\Delta)$ of the Hom space in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ between two supersingular characters. Roughly, one gets that this image is always zero unless the root system is of type $\Phi_{1}\times\Phi_{2}$, where $\Phi_{1}$ is irreducible of rank 2 and $\Phi_{2}$ is a product of $A_{1}$’s, and the two characters are of a very explicit form which we will denote here by $\chi_{1}^{\text{special}}$ and $\chi_{2}^{\text{special}}$, see 5.7 for the details. In order to go from supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ to simple supersingular modules, we then analyse the actions of the group $\Omega$ on these characters and on the Hom spaces between these characters. In particular, given a supersingular character $\chi$ we denote by $\Omega_{\chi}$ its stabiliser in $\Omega$. We also note that $\Omega$ acts on the faces $F$ contained in $\overline{C}$, and we denote the stabiliser of a face $F$ by $\Omega_{F}$. By combining this analysis of the $\Omega$-action on Hom spaces in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$, c.f. 5.20, with the above computation of the image of Hom spaces under ${\operatorname{Ho}}(\Delta)$, we obtain the following:
@@
    Suppose that $\mathbb{G}$ is simply-connected. Then ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ if and only if ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$.
@@
[…]
The above is a combination of 5.15 and 5.24. We then use these results to give a complete classification of the isomorphism classes of simple supersingular modules in the situation where $G$ is a product of $\mathrm{GL}_{n}$’s. Roughly, we then get that two nonisomorphic simple supersingular ${\mathcal{H}}$-modules almost never become isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$ and when they do it is only in a restricted family of cases. In what follows, we denote by $\widetilde{\Omega}$ the lift of $\Omega$ to $\widetilde{W}=N_{G}(T)/(T\cap I)$. This group embeds canonically as a subgroup of ${\mathcal{H}}^{\times}$.
@@
[…]
@@ 1.3. Organisation of the paper
In §3.1-3.2, we study the properties of the functor of parabolic induction in the language of model categories, culminating with the proof of A(i). We only use the fact that parabolic induction is fully faithful, has both a left and a right adjoint, and its left adjoint is exact. Consequently, §3.1 is written in a higher level of generality (c.f. 3.1). In §3.3, we study the relation between the recollement of A(i) and supersingular modules and prove A(ii).
@@ 1.4. Acknowledgments
We would like to thank Jan Kohlhaase and Claudius Heyer for helpful conversations and for giving comments on earlier versions of this paper. The author was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the project Smooth modular representation theory of $p$-adic reductive groups, project number 435414187. The author is also a member of the DFG Research Training Group 2553 ‘Symmetries and Classifying Spaces – Analytic, Arithmetic and Derived’. The financial support of the DFG is gratefully acknowledged.
@@ 1.5. Notation and conventions
A class of objects of a category ${\mathcal{C}}$ will usually be identified with the corresponding full subcategory. If we denote an adjunction by $F:{\mathcal{C}}\rightleftarrows{\mathcal{D}}:U$ then $F$ is always assumed to be left adjoint to $U$. The unit (resp. the counit) of an adjunction will always be denoted by $\eta$ (resp. $\varepsilon$). When such an adjunction is a Quillen adjunction between model categories, we denote by $\mathbf{L}F:{\operatorname{Ho}}({\mathcal{C}})\rightleftarrows{\operatorname{Ho}}({\mathcal{D}}):\mathbf{R}U$ the total derived adjunction between the homotopy categories. Given a functor $F:{\mathcal{C}}\to{\mathcal{D}}$, we denote by ${\operatorname{Im}}(F)$ its essential image. We also say that $F$ preserves (resp. reflects) a property (P) if $*F*$ (resp. ***) has property (P) whenever *** (resp. $*F*$) does.
For any unital ring $R$ we denote by ${\operatorname{Mod}}(R)$ the category of right $R$-modules. Given ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}(R)$ we write ${\operatorname{Hom}}_{R}({\mathfrak{m}},\mathfrak{n})$ for the set of $R$-linear maps from ${\mathfrak{m}}$ to $\mathfrak{n}$, and we denote by ${\operatorname{pd}}_{R}({\mathfrak{m}})$ the projective dimension of ${\mathfrak{m}}$. For any topological group $J$ and field $k$, we denote by ${\operatorname{Rep}}_{k}^{\infty}(J)$ the category of $k$-linear smooth representation of $J$, i.e. the category of all $k$-vector spaces $V$ carrying a $k$-linear action of $J$ such that the stabiliser of any $v\in V$ is open in $J$. The representation $V=k$ equipped with the trivial $J$-action will be denoted by $\mathbf{1}\in{\operatorname{Rep}}_{k}^{\infty}(J)$. Finally, for any $V\in{\operatorname{Rep}}_{k}^{\infty}(J)$ and any subgroup $J^{\prime}\leq J$ we write $V^{J^{\prime}}$ for the vector subspace of $J^{\prime}$-invariant, i.e. of $v\in V$ such that $j\cdot v=v$ for all $j\in J^{\prime}$.

### v4

@@ 1.1. Background and motivation
Let $I$ be a pro-$p$ Iwahori subgroup of $G$ and let ${\mathcal{H}}=k[I\backslash G/I]$ be the corresponding pro-$p$ Iwahori-Hecke algebra of $G$, i.e. the convolution algebra of $I$-biinvariant compactly supported functions on $G$. Write ${\operatorname{Mod}}({\mathcal{H}})$ for the category of right ${\mathcal{H}}$-modules. Then taking $I$-invariants defines a functor $U:{\operatorname{Rep}}_{k}^{\infty}(G)\to{\operatorname{Mod}}({\mathcal{H}})$ which sends non-zero representations to non-zero modules. This functor is expected to provide a strong relation between the two categories, however its behaviour is not sufficiently understood beyond $G=\mathrm{GL}_{2}(\mathbb{Q}_{p})$ and a few related cases (cf. the work of Ollivier [26]). In particular, one of the major open problems in the theory is to determine the relationship between the isomorphism classes of irreducible admissible smooth representations of $G$ and of simple ${\mathcal{H}}$-modules.
In previous work [9] with Kohlhaase, we showed that $U$ restricts to an equivalence between an explicit full subcategory of ${\operatorname{Rep}}_{k}^{\infty}(G)$ and the full subcategory $\mathrm{GProj}({\mathcal{H}})$ of Gorenstein projective ${\mathcal{H}}$-modules. These modules are a key component in the definition of a model structure on ${\operatorname{Mod}}({\mathcal{H}})$ due to Hovey [17]. We used this in [9] to construct and study a model structure on ${\operatorname{Rep}}_{k}^{\infty}(G)$ and the derived functor $\mathbf{R}U$ between the associated homotopy categories. We generally obtained that that the formal properties of this functor are improved at the homotopy level, for instance $\mathbf{R}U$ is essentially surjective. This suggests that one should instead investigate the relationship between the homotopy classes of irreducible admissible smooth representations of $G$ and of simple ${\mathcal{H}}$-modules.
@@
Our main goal in this work is to treat the harder problem of determining whether two nonisomorphic simple ${\mathcal{H}}$-modules of infinite projective dimension can become isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. If $M$ denotes a standard Levi subgroup of $G$ with associated Hecke algebra ${\mathcal{H}}_{M}$ then a key ingredient in Abe’s construction of the simple modules in [1] is the parabolic induction functor ${\operatorname{Ind}}^{{\mathcal{H}}}_{{\mathcal{H}}_{M}}:{\operatorname{Mod}}({\mathcal{H}}_{M})\to{\operatorname{Mod}}({\mathcal{H}})$. We therefore first aim to study the properties of ${\operatorname{Ind}}^{{\mathcal{H}}}_{{\mathcal{H}}_{M}}$ in the language of Hovey’s Gorenstein projective model structure. This allows us to essentially reduce our question to the case of supersingular modules. In many situations of interest, we are then able to obtain a rather complete understanding of the situation (cf. B & C below for details).
@@ 1.2. Main results
    Assume that $k$ is algebraically closed, that the root system of $\mathbb{G}$ is irreducible and that $p\nmid\left|\Omega_{\text{tor}}\right|$. Suppose that ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}})$ has finite length. Then the following are equivalent:
@@
This is proved in 3.13 and 3.32. Roughly speaking, part (i) of A says that the triangulated category ${\operatorname{Ho}}({\mathcal{H}})$ can be decomposed into two orthogonal thick subcategories, one equal to ${\mathcal{S}}^{G}_{\mathcal{M}}$ and the other one equal to ${\operatorname{Ho}}({\mathcal{H}})_{\mathcal{M}}$, where the latter category is embedded fully faithfully into ${\operatorname{Ho}}({\mathcal{H}})$ via either one of the functors $\ell^{G}_{\mathcal{M}}$ or $r^{G}_{\mathcal{M}}$ (cf. 3.9 and the discussion following it for details). We point out that while the proof of (i) is mostly formal, it however relies on the fact that parabolic induction is fully faithful (cf. 2.5), which appears to be a new result.
@@
The proof of part (ii) of A relies on Abe’s classification of the simple ${\mathcal{H}}$-modules in [1], on Koziol’s results on the projective dimension of simple modules from [21], and on a lift to the homotopy level of a computation of Abe of the image of the simple modules under the adjoints of the parabolic induction functor (cf. 3.28). As a consequence of the latter computation, we also obtain a new proof of a result of Koziol and reduce the study of isomorphism classes of simple ${\mathcal{H}}$-modules in ${\operatorname{Ho}}({\mathcal{H}})$ to the case of simple supersingular ${\mathcal{H}}$-modules, cf. 3.29.
Having established this, the next step is to determine when two simple supersingular ${\mathcal{H}}$-modules can become isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. Here the main difficulty is that computing the Hom spaces in ${\operatorname{Ho}}({\mathcal{H}})$ a priori requires to compute explicit Gorenstein projective replacements in ${\operatorname{Mod}}({\mathcal{H}})$. Instead, inspired by the arguments of Koziol in [21] and by the functorial resolutions of Ollivier-Schneider [27], we use Bruhat-Tits theory. If $F$ denotes a face of the Bruhat-Tits building contained in the closure of a chosen chamber $C$, then there is an associated subalgebra ${\mathcal{H}}_{F}$ of ${\mathcal{H}}$. These subalgebras together generate a Gorenstein subalgebra ${\mathcal{H}_{\operatorname{aff}}}$ of ${\mathcal{H}}$ called the affine Hecke algebra. The simple supersingular Hecke modules were classified by Vignéras, cf. [35], and a key input in this classification are the so-called supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$. Our main insight is to consider the ‘diagonal’ functor
@@
which is right Quillen and in fact determines the model structure on ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ via a right transfer property, cf. 5.6. This indicates that we should in principle be able to do computations in the product category $\prod_{F\subseteq\overline{C}}{\operatorname{Ho}}({\mathcal{H}}_{F})$ instead of ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$. The advantage here is that there is no need to compute replacements in order to compute Hom spaces in ${\operatorname{Ho}}({\mathcal{H}}_{F})$ for any $F$, because ${\mathcal{H}}_{F}$ is known to be self-injective and this implies that every ${\mathcal{H}}_{F}$-module is Gorenstein projective.
More specifically, we have that $\Delta$ lifts to a functor ${\operatorname{Ho}}(\Delta):{\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})\to\prod_{F\subseteq\overline{C}}{\operatorname{Ho}}({\mathcal{H}}_{F})$. One major technical input in this paper is the computation of the image under ${\operatorname{Ho}}(\Delta)$ of the Hom space in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ between two supersingular characters. Roughly, one gets that this image is always zero unless the root system is of type $\Phi_{1}\times\Phi_{2}$, where $\Phi_{1}$ is irreducible of rank 2 and $\Phi_{2}$ is a product of $A_{1}$’s, and the two characters are of a very explicit form which we will denote here by $\chi_{1}^{\text{special}}$ and $\chi_{2}^{\text{special}}$, see 5.8 for the details. In order to go from supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ to simple supersingular modules, we then analyse the actions of the group $\Omega$ on these characters and on the Hom spaces between these characters. In particular, given a supersingular character $\chi$ we denote by $\Omega_{\chi}$ its stabiliser in $\Omega$. We also note that $\Omega$ acts on the faces $F$ contained in $\overline{C}$, and we denote the stabiliser of a face $F$ by $\Omega_{F}$. By combining this analysis of the $\Omega$-action on Hom spaces in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$, cf. 5.22, with the above computation of the image of Hom spaces under ${\operatorname{Ho}}(\Delta)$, we obtain the following:
@@
    Suppose that $\mathbb{G}$ is semisimple and simply-connected. Then ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ if and only if ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$.
@@
    Suppose that if one of ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}})$ and ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}}^{\prime})$ contains $\chi_{1}^{\text{special}}$ as a submodule then the other one does not contain $\chi_{2}^{\text{special}}$. Furthermore, if we fix a supersingular character $\chi$ occuring in ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}})$, assume that there exists a face $F\subseteq\overline{C}$ such that $\Omega_{\chi}\subseteq\Omega_{F}$ and ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective. Then ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ if and only if ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$.
The above is a combination of 5.10 and 5.26. We then use these results to give a complete classification of the isomorphism classes of simple supersingular modules in the situation where $G$ is a product of $\mathrm{GL}_{n}$’s. Roughly, we then get that two nonisomorphic simple supersingular ${\mathcal{H}}$-modules almost never become isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$ and when they do it is only in a restricted family of cases. In what follows, we denote by $\widetilde{\Omega}$ the lift of $\Omega$ to $\widetilde{W}=N_{G}(T)/(T\cap I)$. This group embeds canonically as a subgroup of ${\mathcal{H}}^{\times}$.
@@
    If $G$ does have root system of type $A_{2}\times A_{1}\times\cdots\times A_{1}$ (possibly empty product of $A_{1}$’s), then ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ if and only if either ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$ or, up to swapping ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$, we have that $\chi_{1}^{\text{special}}$ and $\chi_{2}^{\text{special}}$ occur as submodules of ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}})$ and ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}}^{\prime})$ respectively and ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ as representations of $\widetilde{\Omega}$.
@@ 1.3. Organisation of the paper
In §3.1-3.2, we study the properties of the functor of parabolic induction in the language of model categories, culminating with the proof of A(i). We only use the fact that parabolic induction is fully faithful, has both a left and a right adjoint, and its left adjoint is exact. Consequently, §3.1 is written in a higher level of generality (cf. 3.1). In §3.3, we study the relation between the recollement of A(i) and supersingular modules and prove A(ii).
@@ 1.4. Acknowledgments
We would like to thank Jan Kohlhaase and Claudius Heyer for helpful conversations and for giving comments on earlier versions of this paper. The author was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the project Smooth modular representation theory of $p$-adic reductive groups, project number 435414187. The author is also a member of the DFG Research Training Group 2553 ‘Symmetries and Classifying Spaces – Analytic, Arithmetic and Derived’. The financial support of the DFG is gratefully acknowledged. We thank the anonymous referees for providing detailed feedback and for making several suggestions which helped improve the paper.
@@ 1.5. Notation and conventions
A class of objects of a category ${\mathcal{C}}$ will usually be identified with the corresponding full subcategory. If we denote an adjunction by $F:{\mathcal{C}}\rightleftarrows{\mathcal{D}}:U$ then $F$ is always assumed to be left adjoint to $U$. The unit (resp. the counit) of an adjunction will always be denoted by $\eta$ (resp. $\varepsilon$). When such an adjunction is a Quillen adjunction between model categories, we denote by $\mathbf{L}F:{\operatorname{Ho}}({\mathcal{C}})\rightleftarrows{\operatorname{Ho}}({\mathcal{D}}):\mathbf{R}U$ the total derived adjunction between the homotopy categories. Given a functor $F:{\mathcal{C}}\to{\mathcal{D}}$, we denote by ${\operatorname{Im}}(F)$ its essential image. We also say that $F$ preserves (resp. reflects) a property (P) if $F*$ (resp. *) has property (P) whenever * (resp. $F*$) does.
For any unital ring $R$ we denote by ${\operatorname{Mod}}(R)$ the category of right $R$-modules. Given ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}(R)$ we write ${\operatorname{Hom}}_{R}({\mathfrak{m}},\mathfrak{n})$ for the set of $R$-linear maps from ${\mathfrak{m}}$ to $\mathfrak{n}$, and we denote by ${\operatorname{pd}}_{R}({\mathfrak{m}})$ and ${\operatorname{id}}_{R}({\mathfrak{m}})$ the projective and injective dimensions of ${\mathfrak{m}}$, respectively. For any topological group $J$ and field $k$, we denote by ${\operatorname{Rep}}_{k}^{\infty}(J)$ the category of $k$-linear smooth representations of $J$, i.e. the category of all $k$-vector spaces $V$ carrying a $k$-linear action of $J$ such that the stabiliser of any $v\in V$ is open in $J$. The representation $V=k$ equipped with the trivial $J$-action will be denoted by $\mathbf{1}\in{\operatorname{Rep}}_{k}^{\infty}(J)$. Finally, for any $V\in{\operatorname{Rep}}_{k}^{\infty}(J)$ and any subgroup $J^{\prime}\leq J$ we write $V^{J^{\prime}}$ for the vector subspace of $J^{\prime}$-invariants, i.e. of $v\in V$ such that $j\cdot v=v$ for all $j\in J^{\prime}$.